OpenStudy (anonymous):

*fan and medal awarded* For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 -x - 2y = -10 -2y + 10 2y + 10 2y - 10 -2y - 10

OpenStudy (anonymous):

how do u add x

OpenStudy (anonymous):

what is x. x isnt a number so i cant add it. please explain.

OpenStudy (hba):

-x - 2y = -10 Shift x to the other side of the '= sign'

OpenStudy (anonymous):

okay after adding x what would iget cuz your not really helping. i can do the add 10 thing

OpenStudy (hba):

Ok add 10 :)

OpenStudy (anonymous):

what do i get after adding x.

OpenStudy (anonymous):

You would try to isolate the x term: -x - 2y = -10 First move the -2y over +2y +2y ---------------- -x = 2y - 10 Now get rid of the negative on the x by dividing by -1 ----------- -1 x= -2y + 10 <-- this is what you would plug in